OpenStudy (anonymous):

Sandy invested $2,500 in a corporate bond that pays 9% interest compounded continuously. How long will it take for Sandy's investment to triple? (Reminder: The formula for interest compounded continuously is , where A is the amount in the account, P is the principal, r is the interest rate, and t is the time in years.) What equation correctly shows the information from the problem substituted into the equation?

OpenStudy (anonymous):

Do you know the formula? A = Pe^(rt)

OpenStudy (anonymous):

^^ That's the correct formula. But there's 2 formulas for compound interest, don't use the above formula for every equation.

OpenStudy (anonymous):

A= Pe^(rt) is for equations dealing with compounding continuously

OpenStudy (anonymous):

2,500=Pe^(9 3) ?

OpenStudy (anonymous):

^^ yep, there's also a formula for dealing compound monthly, weekly, etc. Just didn't want OP to think there's only one formula

OpenStudy (anonymous):

No, 2500(3) = 2500e^(.009t)

OpenStudy (anonymous):

P is the initial amount Sandy puts in, which is 2500. What the equation is equal to is the total amount AFTER the investment is down being compounded

OpenStudy (anonymous):

O.k got it. So if I were to divide both sides of the equation by 2,500... The result would be... 3=e^(.09t)?

OpenStudy (anonymous):

Yes, then you use one of the rules of logs to take out the e and solve for t

OpenStudy (anonymous):

Basically rewrite the exponential equation (stated above) in logarithmic form?

OpenStudy (anonymous):

I suggest you use natural log.

OpenStudy (anonymous):

Yeah, what Denebel said, so it becomes, ln3 = .09t If denebel would kindly check for me, I haven't dealt with logs, lns etc. in years

OpenStudy (anonymous):

Denebel? Is this correct? I would really appreciate your confirmation :)

OpenStudy (anonymous):

It's correct, I solved it and worked it out, and checked, the answer you get works

OpenStudy (anonymous):

Yes